Circular No. 6565

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GX 17+2 M. van der Klis, J. Homan, and R. Wijnands, Astronomical Insti-

tute ‘Anton Pannekoek’, University of Amsterdam (UA); E. Kuulkers, Oxford University; F. K. Lamb and D. Psaltis, University of Illinois at Urbana/Champaign; S. Dieters, University of Alabama at Huntsville (UAH); J. van Paradijs, UAH and UA; W. H. G. Lewin, Massachusetts Institute of Technology; and B. Vaughan, California Institute of Technology, report: “RXTE observations of the Z-source and low-mass-x-ray-binary GX 17+2 that we obtained during Feb. 6–9 show quasiperiodic oscillations (QPOs) at the kHz level. We find two peaks in the power spectra of the x-ray countrate variations. The frequency of the lower-frequency peak decreases from ∼ 880 to 682 ± 5 Hz when the countrate increases from 1600 to 1920 counts/s. This is consistent with a positive correlation between frequency and mass-accretion rate, as the source is on the normal branch in the x-ray color-color diagram during our observations, where countrate and accretion rate are thought to be anticorrelated. The higher-frequency peak is detected only at high countrate, at a frequency of 988 ± 3 Hz. The peak separation is then 306 ± 5 Hz. Peak widths range from 30 to 200 Hz, and QPO amplitudes are between 3 and 5 percent (rms), with no significant dependence on countrate. We also observe simultaneous horizontal-branch oscillations of ∼ 60 Hz, but no 6-Hz QPOs, consistent with the source being in the upper-normal branch. The kHz QPOs were detected in the band 8– 60 keV and are very difficult to see in the more usual 2–60-keV band. X-ray observers are advised to use high time-resolution modes, which allow isolating the high-energy photons, in ongoing observational programs of bright low-mass x-ray binaries in order to be able to detect the phenomenon.”

COMET 46P/WIRTANEN J.-L. Bertaux, on behalf of the SWAN team at Service d’Aeronomie,

Paris, and the Finnish Meteorological Institute, Helsinki, writes: “The Lyman-α emission of the hydrogen envelope of comet 46P was observed for the first time on Feb. 10 (at r = 1.15 AU) with the SWAN experiment onboard the SOHO spacecraft at the L_{1} Lagrange point (the spacecraft being 1.61 AU from the comet). From the peak intensity of 44 rayleighs in SWAN’s 1^{o} field-of-view, the water-production rate is estimated at 7 × 10^{27} molecules/s, using a comparison with a simple Haser model.”

1997 February 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green